# On De Giorgi’s lemma for variational interpolants

### v2

@@ Abstract
Variational interpolants are an indispensable tool for the construction of gradient-flow solutions via the Minimizing Movement Scheme. De Giorgi’s lemma provides the associated discrete energy-dissipation inequality. It was originally developed for metric gradient systems. Drawing from this theory we study the case of generalized gradient systems in Banach spaces, where a refined theory allows us to extend the validity of the discrete energy-dissipation inequality and to establish it as an equality. For the latter we have to impose the condition of radial differentiability of the dissipation potential. Several examples are discussed to show how sharp the results are.
@@ 1 Introduction
The Minimizing Movement Scheme (MMS) was introduced by De Giorgi in [De93] for constructing solutions for gradient flows in abstract spaces. Since then, the MMS has developed into a versatile tool for analyzing gradient systems in Hilbert spaces, Banach space, and metric spaces. In this paper, we address the specific tool called “variational interpolant”, also called “De Giorgi’s interpolant” that was first introduced in [Amb95, Lem. 2.5] and further developed in [AGS05]. A generalization to the Banach spaces was done in [MRS13a, Lem. 6.1]. Variational interpolants generalize the idea of piecewise affine interpolants in linear spaces, or geodesic interpolants in geodesic spaces, such that they are applicable in more general situations, namely in general metric spaces. However, even in the cases of geodesic spaces, including Banach and Hilbert spaces, they are useful if the energy functional is not geodesically semiconvex. In general, variational interpolants are no longer continuous in time and hence, the desired discrete energy-dissipation estimate is more difficult to obtain. It is exactly this estimate, which is established in the so-called “De Giorgi’s lemma”. The purpose of this paper is twofold: (i) we generalize the validity of the lemma in the Banach setup and (ii) we discuss the question why and when the discrete energy-dissipation estimate is an equality.
@@
To be more precise, we now introduce our approach in more detail by comparing the theory in metric spaces $(M,{\mathcal{D}})$ and in Banach spaces $(X;\|\cdot\|)$ in parallel. Following [RMS08] (see also [Mie23]) we consider a generalized metric gradient system $(M,{\mathcal{E}},{\mathcal{D}},\psi)$, subsequently abbreviated by $\mathrm{gMGS}$; see Definition 2.1 for the precise definition. For a given initial value $u^{o}\in M$ with ${\mathcal{E}}(u^{o})<\infty$, it is the aim to construct a curve $u:{[0,\infty[}\to M$ of maximal slope emanating from $u^{o}$, i.e. $u$ must satisfy for all $t>0$
${\mathcal{E}}(u(t))+\int_{0}^{t}\Big{(}\psi\big{(}|u^{\prime}|(s)\big{)}+\psi^{*}\big{(}|\partial{\mathcal{E}}|(u(s))\big{)}\Big{)}\;\!\mathrm{d}s={\mathcal{E}}(u(0))\quad\text{and}\quad u(0)=u^{o},$ (1.1a)
@@
where $|u^{\prime}|\geq 0$ denotes the metric speed of $u$ and $|\partial{\mathcal{E}}|(u)\geq 0$ denotes the metric slope, see [AGS05]. The case of general dissipation functions $\psi:{[0,\infty[}\to{[0,\infty[}$ (lower semicontinuous, convex, $\psi(0)=0)$, and superlinear) was introduced in [RMS08], the choice $\psi(r)=\frac{1}{2}r^{2}$ gives the classical notion of curve of maximal slope of [Amb95], while $\phi(r)=\frac{1}{p}r^{p}$ leads to $p$-curves of maximal slopes as in [AGS05].
@@
[…]
where $u^{\prime}$ is the distributional derivative of $u\in\mathrm{AC}([0,T];X)$, and $\xi:{]0,\infty[}\to X^{*}$ is a selection in the multivalued Fréchet subdifferential ${\mathrm{D}}{\mathcal{E}}(u)\subset X^{*}$ of ${\mathcal{E}}$, see Section 3.1.
With an initial value $u^{o}\in M$ and a time step $\tau>0$ the metric and the Banach-space MMS are defined via $u^{0}_{\tau}=u^{o}$ and
@@
De Giorgi’s lemma, which was first published in [Amb95, Lem. 2.5], now provides a discrete counterpart to the energy-dissipation balances in (1.1a) and (1.1b), namely for all $\sigma\in{]0,\tau]}$ we have the so-called De Giorgi’s estimates
@@
${\mathcal{E}}(\widetilde{u}^{\tau}(k\tau{+}\sigma)\big{)}+\sigma\,{\mathcal{R}}\big{(}\tfrac{\displaystyle 1}{\displaystyle\sigma}(\widetilde{u}^{\tau}(k\tau{+}\sigma){-}u^{k}_{\tau})\big{)}+\int_{0}^{\sigma}{\mathcal{R}}^{*}\big{(}{-}\xi(k\tau{+}\rho))\big{)}\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{k}_{\tau})$ (1.4b)
[…]
@@
For general differentiable dissipation functions $\psi$, De Giorgi’s estimate (1.4a) was first established in [RMS08, Lem. 4.5], while $\psi(r)=\frac{1}{p}r^{p}$ is treated in [AGS05]. The Banach-space case (1.4b) appears first in [MRS13a, Lem. 6.1], but the result therein relies on the condition of radial differentiability of ${\mathcal{R}}$, namely
@@
So far, our general overview and introduction shows a complete analogy between the metric case and the Banach-space setting. Even the condition of radial differentiability of ${\mathcal{R}}$ corresponds to the condition of differentiability of $\psi$. However, the methods for establishing the so-called De Giorgi’s estimates (1.4a) and (1.4b) involve quite different techniques. In particular, for $\mathrm{gBGS}$ we can exploit the linear structure of $X$ and thus obtain an Euler-Lagrange equation for the minimizers $\widetilde{u}_{\sigma}:=\widetilde{u}^{\tau}(k\tau{+}\sigma)$ (keep $k$ fixed, w.l.o.g. $k=0$), namely
@@
This observation (which we shall revisit in Section 4.4), motivates our first main result, see Theorem 4.1, that De Giorgi’s estimate (1.4b) is indeed an equality, then called De Giorgi’s identity. This result is established for all measurable variational interpolants, under the sole additional assumption of radial differentiability of ${\mathcal{R}}$, cf. (1.5). However, for this version of De Giorgi’s identity, it is essential to be more specific with the choice of $\widetilde{\xi}_{\sigma}\in{\mathrm{D}}{\mathcal{E}}(\widetilde{u}_{\sigma})$: one has to restrict to those $\xi\in{\mathrm{D}}{\mathcal{E}}(\widetilde{u}_{\sigma})$ that minimize ${\mathcal{R}}^{*}({-}\xi)$ subject to the constraint of satisfying the Euler-Lagrange equation (1.6), see (3.14) for the precise definition. We refer to Example 3.6 for a very simple case, where this restriction is essential for the validity of De Giorgi’s estimate as an identity.
[…]
@@
$\mathsf{d}^{-}_{\rho}(u^{o})$ $=\inf\big{\{}\,{\mathcal{D}}(u^{o},u)\>\big{|}\>u\in\mathrm{Argmin}\big{(}{\mathcal{D}}(u^{o},\cdot)^{2}/(2\rho)+{\mathcal{E}}(\cdot)\big{)}\,\big{\}},$
$\mathsf{d}^{+}_{\rho}(u^{o})$ $=\sup\big{\{}\,{\mathcal{D}}(u^{o},u)\>\big{|}\>u\in\mathrm{Argmin}\big{(}{\mathcal{D}}(u^{o},\cdot)^{2}/(2\rho)+{\mathcal{E}}(\cdot)\big{)}\,\big{\}}.$
@@
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma\,\psi\big{(}\tfrac{\displaystyle 1}{\displaystyle\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}+\int_{\rho=0}^{\sigma}\psi^{*}\Big{(}\psi^{\prime}\big{(}\tfrac{\displaystyle 1}{\displaystyle\rho}\mathsf{d}^{\pm}_{\rho}(u^{o})\big{)}\Big{)}\;\!\mathrm{d}\rho={\mathcal{E}}(u^{o}).$ (1.7)
@@
Clearly, the metric De Giorgi’s estimate (1.4a) follows easily from (1.7) by inserting the slope inequality
$|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big{(}\tfrac{\displaystyle 1}{\displaystyle\sigma}{\mathcal{D}}(\widetilde{u}_{\sigma},u^{k}_{\tau})\big{)}\,\big{|}\partial({-}{\mathcal{D}}(u^{o},\cdot)\big{|}(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big{(}\tfrac{\displaystyle 1}{\displaystyle\sigma}{\mathcal{D}}(\widetilde{u}_{\sigma},u^{k}_{\tau})\big{)},$ (1.8)
see Proposition 2.9. The latter can be seen as a metric counterpart of the Euler-Lagrange equation. If one of the two inequalities in (1.8) is strict, then De Giorgi’s identity is lost. The last inequality is strict, if ${\mathcal{D}}$ is not a length distance, which means that the metric De Giorgi estimate can only hold in geodesic spaces. However, even there the first inequality may be strict. In Theorem 2.14 we show that full continuity and a uniform slope estimate are sufficient to establish De Giorgi’s identity in geodesic metric spaces.
@@ 2 The metric case
    ${\mathcal{E}}:M\to(-\infty,\infty]$ is a proper (i.e. with non-empty domain $\mathop{\mathrm{dom}}({\mathcal{E}})$) lsc functional;
@@
Nonetheless, we have opted for omitting the role of $\mathscr{T}$ in the discussion of the metric case in order to avoid overburdening it, on the one hand, and to highlight the purely metric flavour of the arguments, on the other hand. Instead, in the Banach setup it will be convenient to encompass the weak topology in the picture.
In the setup of a $\mathrm{gMGS}$ $(M,{\mathcal{E}},{\mathcal{D}},\psi)$, the classical notion of curve of maximal slope is extended by the following definition (cf. [Mie23, Def. 4.8]). To simplify the arguments, we fix an arbitrary $T>0$ and confine the discussion to evolutions on the compact time interval $[0,T]$.
@@
In [RMS08, Prob. 2.6] an alternative definition for the above concept was given, claiming for the curve $u\in\mathrm{AC}([0,T];M)$ the pointwise estimate
@@
The Minimizing Movement Scheme for constructing curves of maximal slope fulfilling the initial condition $u(0)=u_{0}$ for an assigned initial datum $u_{0}\in M$ then reads as follows: given a time step $\tau>0$ inducing a (uniform, without loss of generality) partition $\mathscr{P}_{\tau}=\{t_{\tau}^{k}\}_{k=1}^{K_{\tau}}$ of the interval $[0,T]$, starting from $u_{\tau}^{0}:=u_{0}$, find $(u^{k}_{\tau})_{k=1}^{K_{\tau}}$ such that
@@
$\mathrm{Min}_{u\in M}\Phi_{\sigma}(u^{o};u)\quad\text{with }\ \Phi_{\sigma}(u^{o};u):=\sigma\psi\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},u)\big{)}+{\mathcal{E}}(u),\quad\sigma>0,$ (2.4)
for a fixed $u^{o}\in M$. We will also use the following notation
$\phi(u^{o};\sigma)=\inf\big{\{}\,\Phi_{\sigma}(u^{o};u)\>\big{|}\>u\in M\,\big{\}}\ \text{ and }J_{\sigma}(u^{o})=\mathrm{ArgMin}\big{\{}\,\Phi_{\sigma}(u^{o};u)\>\big{|}\>u\in M\,\big{\}}.$ (2.5)
@@
$\mathsf{d}^{-}_{\sigma}(u^{o}):=\inf\big{\{}\,{\mathcal{D}}(u^{o},u)\>\big{|}\>u\in J_{\sigma}(u^{o})\,\big{\}}\quad\text{and}\quad\mathsf{d}^{+}_{\sigma}(u^{o}):=\sup\big{\{}\,{\mathcal{D}}(u^{o},u)\>\big{|}\>u\in J_{\sigma}(u^{o})\,\big{\}}.$
@@
    $J_{\sigma}(u^{o})\neq\emptyset\text{ for all }\sigma\in(0,\sigma_{*})\ \text{ and all }u^{o}\in\mathop{\mathrm{dom}}({\mathcal{E}}).$ (2.6)
@@
Whenever the generalized metric gradient system is individuated by a quintuple $(M,\mathscr{T},{\mathcal{E}},{\mathcal{D}},\psi)$ such that the topology $\mathscr{T}$ is compatible with ${\mathcal{D}}$, (2.6) follows by a coercivity property of this type: there exists $\sigma_{*}>0$ such that for all $\sigma\in(0,\sigma_{*})$ and all $u^{o}\in M$, for any sequence $(u_{n})_{n}\subset M$
[…]
@@
$\text{there exists a \emph{measurable} selection }(0,\sigma_{*})\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o}).$ (2.8)
To show this, let us consider the multi-valued mapping $\Gamma:(0,\infty)\rightrightarrows X$, $\Gamma(\sigma):=J_{\sigma}(u^{o})$. It is easy to check that $\Gamma$ is upper semicontinuous from ${\mathbb{R}}$ to $(M,{\mathcal{D}})$ in that it fulfills for every $(\sigma_{n})_{n},\,\sigma\in(0,\infty)$
$\sigma_{n}\to\sigma\ \Longrightarrow\ \mathop{\mathrm{Ls}}_{n\to\infty}\Gamma(\sigma_{n})\subset\Gamma(\sigma)\,,$ (2.9a)
where the Kuratowski upper limit (cf., e.g., [AmT04, Def. 4.4.13]) of the sequence of closed sets $(\Gamma(\sigma_{n}))_{n}$ is defined by
$u\in\mathop{\mathrm{Ls}}_{n\to\infty}\Gamma(\sigma_{n})\ \Longleftrightarrow\ \exists\,(\sigma_{n_{k}})_{k},\,(u_{k})_{k}\text{ with }\begin{cases}\displaystyle u_{k}\in\Gamma(\sigma_{n_{k}})\text{ for all }k\in{\mathbb{N}}\,,\\ \lim_{k\to\infty}{\mathcal{D}}(u,u_{k})=0\,.\end{cases}$ (2.9b)
Since $\Gamma$ is upper semicontinuous, its graph is a Borel subset of $(0,\infty)\times X$, and the von Neumann-Aumann selection theorem [CaV77, Thm. 3.22] applies, yielding (2.8).
@@
Definition 2.7 (De Giorgi’s variational interpolant)
@@
The cornerstone of the proof that, as $\tau\downarrow 0$, (a subsequence of) the sequence $(\widetilde{u}_{\tau})_{\tau}$ converges to a curve of maximal slope is the discrete estimate obtained by applying De Giorgi’s estimate (1.4a) to the interpolant $\widetilde{u}_{\tau}$. The next section revolves around the validity of (1.4a) as an equality.
2.1 Metric energy identity and De Giorgi’s estimate: statements and examples

### v3

@@ Abstract
Variational interpolants are an indispensable tool for the construction of gradient-flow solutions via the Minimizing Movement Scheme. The De Giorgi lemma provides the associated discrete energy-dissipation inequality. It was originally developed for metric gradient systems. Drawing from this theory we study the case of generalized gradient systems in Banach spaces, where a refined theory allows us to extend the validity of the discrete energy-dissipation inequality and to establish it as an equality. For the latter we have to impose the condition of radial differentiability of the dissipation potential. Several examples are discussed to show how sharp the results are.
@@ 1 Introduction
The Minimizing Movement Scheme (MMS) was introduced by De Giorgi in [De93] for constructing solutions for gradient flows in abstract spaces. Since then, the MMS has developed into a versatile tool for analyzing gradient systems in Hilbert spaces, Banach space, and metric spaces. In this paper, we address the specific tool called “variational interpolant”, also called “De Giorgi interpolant” that was first introduced in [Amb95, Lem. 2.5] and further developed in [AGS05]. A generalization to the Banach spaces was done in [MRS13, Lem. 6.1]. Variational interpolants generalize the idea of piecewise affine interpolants in linear spaces, or geodesic interpolants in geodesic spaces, in such a way that they turn out to be applicable in more general situations, namely in general metric spaces. However, even in the cases of geodesic spaces, including Banach and Hilbert spaces, they are useful if the energy functional is not geodesically semi-convex. In general, variational interpolants are no longer continuous in time and hence, the desired discrete energy-dissipation estimate is more difficult to obtain. It is exactly this estimate, which is established in the so-called “De Giorgi lemma”. The purpose of this paper is twofold: (i) we generalize the validity of the lemma in the Banach setup and (ii) we discuss the question why and when the discrete energy-dissipation estimate is an equality.
To be more precise, we now introduce our approach in more detail by comparing the theory in metric spaces $(M,{\mathcal{D}})$ and in Banach spaces $(X;\|\cdot\|)$ in parallel. It has to be emphasized that the Banach setup we are going to address is not contained in the metric one. Indeed, while in the latter we shall essentially confine the discussion to the case in which the dissipation mechanism is encoded by the interplay of the metric $\mathcal{D}$} with a convex scalar function $\psi:{[0,\infty[}\to{[0,\infty[}$, in the former we will address general dissipation potentials ${\mathcal{R}}:X\to[0,\infty]$, for instance ${\mathcal{R}}(v)=\int_{\Omega}(\frac{1}{2}v^{2}{+}\frac{1}{4}v^{4})\;\!\mathrm{d}x$ for $X={\mathrm{L}}^{4}(\Omega)$.
@@
Following [RMS08] (see also [Mie23]) we consider a generalized metric gradient system $(M,{\mathcal{E}},{\mathcal{D}},\psi)$, subsequently abbreviated by $\mathrm{gMGS}$; see Definition 2.1 for the precise definition. For a given initial value $u^{\circ}\in M$ with ${\mathcal{E}}(u^{\circ})<\infty$, it is the aim to construct a curve $u:{[0,\infty[}\to M$ of maximal slope emanating from $u^{\circ}$, i.e. $u$ must satisfy for all $t>0$
${\mathcal{E}}(u(t))+\int_{0}^{t}\Big(\psi\big(|u^{\prime}|(s)\big)+\psi^{*}\big(|\partial{\mathcal{E}}|(u(s))\big)\Big)\;\!\mathrm{d}s={\mathcal{E}}(u(0))\quad\text{and}\quad u(0)=u^{\circ},$ (1.1a)
@@
where $|u^{\prime}|\geq 0$ denotes the metric speed of $u$ and $|\partial{\mathcal{E}}|(u)\geq 0$ denotes the metric slope, see [AGS05], and $\psi^{*}$ is the convex conjugate of $\psi$. The case of general dissipation functions $\psi:{[0,\infty[}\to{[0,\infty[}$ (lower semicontinuous, convex, $\psi(0)=0)$, and superlinear) was introduced in [RMS08], the choice $\psi(r)=\frac{1}{2}r^{2}$ gives the classical notion of curve of maximal slope of [Amb95], while $\phi(r)=\frac{1}{p}r^{p}$ leads to $p$-curves of maximal slopes as in [AGS05].
@@
${\mathcal{E}}(u(t))+\int_{0}^{t}\!\Big({\mathcal{R}}\big(u^{\prime}(s)\big)+{\mathcal{R}}^{*}({-}\xi(s))\Big)\;\!\mathrm{d}s={\mathcal{E}}(u(0))$ $\text{ for all }t>0,$ (1.1b)
$\text{ and }\xi(s)\in{\mathrm{D}}{\mathcal{E}}(u(s))$ $\text{ for a.a.\ }s\geq 0,$
where $u^{\prime}$ is the distributional derivative of $u\in\mathrm{AC}([0,T];X)$, and the convex conjugate ${\mathcal{R}}^{*}:X^{*}\to[0,\infty)$ is evaluated along a selection $\xi:{]0,\infty[}\to X^{*}$ in the multivalued Fréchet subdifferential ${\mathrm{D}}{\mathcal{E}}(u)\subset X^{*}$ of $\mathcal{E}$}, see Section 3.1.
With an initial value $u^{\circ}\in M$ and a time step $\tau>0$ the metric and the Banach-space MMS are defined via $u^{0}_{\tau}=u^{\circ}$ and
@@
The De Giorgi lemma, which was first published in [Amb95, Lem. 2.5], provides a discrete counterpart to the energy-dissipation balances in (1.1a) and (1.1b), namely for all $\sigma\in{]0,\tau]}$ we have the so-called De Giorgi estimates
@@
${\mathcal{E}}(\widetilde{u}^{\tau}(k\tau{+}\sigma)\big)+\sigma\,{\mathcal{R}}\big(\tfrac{\displaystyle 1}{\displaystyle\sigma}(\widetilde{u}^{\tau}(k\tau{+}\sigma){-}u^{k}_{\tau})\big)+\int_{0}^{\sigma}{\mathcal{R}}^{*}\big({-}\xi(k\tau{+}\rho)\big)\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{k}_{\tau})$ (1.4b)
 $\hskip 180.00027pt\text{ for some }\xi(t)\in{\mathrm{D}}{\mathcal{E}}(\widetilde{u}^{\tau}(t))\text{ for a.a.\ }t>0.$
@@
For general differentiable dissipation functions $\psi$, the De Giorgi estimate (1.4a) was first established in [RMS08, Lem. 4.5], while $\psi(r)=\frac{1}{p}r^{p}$ is treated in [AGS05]. The Banach-space case (1.4b) appears first in [MRS13, Lem. 6.1], but the result therein relies on the condition of radial differentiability of $\mathcal{R}$}, namely
@@
So far, our general overview and introduction shows a complete analogy between the metric case and the Banach-space setting. Even the condition of radial differentiability of $\mathcal{R}$} corresponds to the condition of differentiability of $\psi$. However, the methods for establishing the so-called De Giorgi estimates (1.4a) and (1.4b) involve quite different techniques. In particular, for $\mathrm{gBGS}$ we can exploit the linear structure of $X$ and thus obtain an Euler-Lagrange equation for the minimizers $\widetilde{u}_{\sigma}:=\widetilde{u}^{\tau}(k\tau{+}\sigma)$ (keep $k$ fixed, w.l.o.g. $k=0$), namely
@@
This observation (which we shall revisit in Section 4.5), motivates our first main result, see Theorem 4.2, that the De Giorgi estimate (1.4b) is indeed an equality, then called De Giorgi identity. This result is established for all measurable variational interpolants, under the sole additional assumption of radial differentiability of $\mathcal{R}$}, cf. (1.5). However, for this version of the De Giorgi identity, it is essential to be more specific with the choice of $\widetilde{\xi}_{\sigma}\in{\mathrm{D}}{\mathcal{E}}(\widetilde{u}_{\sigma})$: one has to restrict to those $\xi\in{\mathrm{D}}{\mathcal{E}}(\widetilde{u}_{\sigma})$ that minimize ${\mathcal{R}}^{*}({-}\xi)$ subject to the constraint of satisfying the Euler-Lagrange equation (1.6), see (3.17) for the precise definition. We refer to Example 3.9 for a very simple case, where this restriction is essential for the validity of the De Giorgi estimate as an identity.
For the case of a general $\mathcal{R}$}, dropping radial differentiability, we are able to establish the De Giorgi estimate (1.4b) if $X$ is a reflexive Banach space. In fact, along with (1.4b) we shall also obtain a refined estimate, involving a force selection $\widetilde{\xi}_{\sigma}\in{\mathrm{D}}{\mathcal{E}}(\widetilde{u}_{\sigma})$ that also satisfies the Euler-Lagrange equation (1.6). In fact, we shall refer to (1.4b) as the simple De Giorgi estimate, which will be enhanced to the improved De Giorgi estimate keeping track of (1.6). Both, the simple and the improved De Giorgi estimates will be proved in our second main result, Theorem 4.12. For this, we use a Yosida-Moreau regularization ${\mathcal{R}}_{\eta}$ of $\mathcal{R}$} with an equivalent norm $\|\cdot\|$ such that $u\mapsto\|u\|^{2}$ is differentiable. Then, ${\mathcal{R}}_{\eta}$ is differentiable, in particular also radially differentiable, and for the corresponding $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}}_{\eta})$ the De Giorgi identity holds thanks to Theorem 4.2. It can be shown that in the limit passage $\eta\to 0^{+}$ the De Giorgi estimate survives.
@@
$\mathsf{d}^{-}_{\rho}(u^{\circ})$ $=\inf\big\{\,{\mathcal{D}}(u^{\circ},u)\>\big|\>u\in\mathrm{Argmin}\big({\mathcal{D}}(u^{\circ},\cdot)^{2}/(2\rho)+{\mathcal{E}}(\cdot)\big)\,\big\},$
$\mathsf{d}^{+}_{\rho}(u^{\circ})$ $=\sup\big\{\,{\mathcal{D}}(u^{\circ},u)\>\big|\>u\in\mathrm{Argmin}\big({\mathcal{D}}(u^{\circ},\cdot)^{2}/(2\rho)+{\mathcal{E}}(\cdot)\big)\,\big\}.$
@@
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma\,\psi\big(\tfrac{\displaystyle 1}{\displaystyle\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)+\int_{\rho=0}^{\sigma}\psi^{*}\Big(\psi^{\prime}\big(\tfrac{\displaystyle 1}{\displaystyle\rho}\mathsf{d}^{\pm}_{\rho}(u^{\circ})\big)\Big)\;\!\mathrm{d}\rho={\mathcal{E}}(u^{\circ}).$ (1.7)
@@
Clearly, the metric De Giorgi estimate (1.4a) follows easily from (1.7) by inserting the slope inequality
$|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big(\tfrac{\displaystyle 1}{\displaystyle\sigma}{\mathcal{D}}(\widetilde{u}_{\sigma},u^{k}_{\tau})\big)\,\big|\partial({-}{\mathcal{D}}(u^{\circ},\cdot)\big|(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big(\tfrac{\displaystyle 1}{\displaystyle\sigma}{\mathcal{D}}(\widetilde{u}_{\sigma},u^{k}_{\tau})\big),$ (1.8)
see Proposition 2.9. The latter can be seen as a metric counterpart of the Euler-Lagrange equation. If one of the two inequalities in (1.8) is strict, then the De Giorgi identity is lost. The last inequality is strict, if $\mathcal{D}$} is not a length distance, which means that the metric De Giorgi estimate can only hold in geodesic spaces. However, even there the first inequality may be strict. In Theorem 2.15 we show that full continuity and a uniform slope estimate are sufficient to establish the De Giorgi identity in geodesic metric spaces.
@@ 2 The metric case
    ${\mathcal{E}}:M\to(-\infty,\infty]$ is a proper (i.e., with non-empty domain $\mathop{\mathrm{dom}}({\mathcal{E}})$) lsc functional;
@@
Nonetheless, we have opted for omitting the role of $\mathscr{T}$ in the discussion of the metric case in order to avoid overburdening it, on the one hand, and to highlight the purely metric flavor of the arguments, on the other hand. Instead, in the Banach setup it will be convenient to encompass the weak topology in the picture.
In the setup of a $\mathrm{gMGS}$ $(M,{\mathcal{E}},{\mathcal{D}},\psi)$, the classical notion of curve of maximal slope is extended by the following definition (cf. [Mie23, Def. 4.8]), which brings into play the convex conjugate $\psi^{*}:[0,\infty[\to[0,\infty[$, $\psi^{*}(s)=\sup_{r\in[0,\infty[}(sr{-}\psi(r))$, of $\psi$. To simplify the arguments, we fix an arbitrary $T>0$ and confine the discussion to evolutions on the compact time interval $[0,T]$.
@@
In [RMS08, Prob. 2.6] an alternative definition for the above concept was given, imposing for the curve $u\in\mathrm{AC}([0,T];M)$ the pointwise estimate
@@
The Minimizing Movement Scheme for constructing curves of maximal slope fulfilling the initial condition $u(0)=u_{0}$, for an assigned initial datum $u_{0}\in M$, then reads as follows: given a time step $\tau>0$, inducing a (uniform, without loss of generality) partition $\mathscr{P}_{\tau}=\{t_{\tau}^{k}\}_{k=1}^{K_{\tau}}$ of the interval $[0,T]$, starting from $u_{\tau}^{0}:=u_{0}$ find $(u^{k}_{\tau})_{k=1}^{K_{\tau}}$ such that
@@
$\mathop{\mathrm{Min}}_{u\in M}\Phi_{\sigma}(u^{\circ};u)\quad\text{with }\ \Phi_{\sigma}(u^{\circ};u):=\sigma\psi\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},u)\big)+{\mathcal{E}}(u),\quad\sigma>0,$ (2.4)
for a fixed $u^{\circ}\in M$. We will also use the following notation
$\phi(u^{\circ};\sigma)=\inf\big\{\,\Phi_{\sigma}(u^{\circ};u)\>\big|\>u\in M\,\big\}\ \text{ and }J_{\sigma}(u^{\circ})=\mathrm{Argmin}\big\{\,\Phi_{\sigma}(u^{\circ};u)\>\big|\>u\in M\,\big\}.$ (2.5)
@@
$\mathsf{d}^{-}_{\sigma}(u^{\circ}):=\inf\big\{\,{\mathcal{D}}(u^{\circ},u)\>\big|\>u\in J_{\sigma}(u^{\circ})\,\big\}\quad\text{and}\quad\mathsf{d}^{+}_{\sigma}(u^{\circ}):=\sup\big\{\,{\mathcal{D}}(u^{\circ},u)\>\big|\>u\in J_{\sigma}(u^{\circ})\,\big\}.$
@@
    $J_{\sigma}(u^{\circ})\neq\emptyset\text{ for all }\sigma\in(0,\sigma_{*})\ \text{ and all }u^{\circ}\in\mathop{\mathrm{dom}}({\mathcal{E}}).$ (2.6)
@@
Whenever the generalized metric gradient system is individuated by a quintuple $(M,\mathscr{T},{\mathcal{E}},{\mathcal{D}},\psi)$ such that the topology $\mathscr{T}$ is compatible with $\mathcal{D}$}, (2.6) follows from a coercivity property of the following type: there exists $\sigma_{*}>0$ such that for all $\sigma\in(0,\sigma_{*})$ and all $u^{\circ}\in M$, for all sequences $(u_{n})_{n}\subset M$ we have the implication
$\sup_{n}\Phi_{\sigma}(u^{\circ};u_{n})<+\infty\ \Longrightarrow\ (u_{n})_{n}\text{ admits a $\mathscr{T}$-converging subsequence}.$ (2.7)
@@
Since this paper is focused on the one-step minimum problem (2.4) and its ‘starting’ point $u^{\circ}$ is fixed, throughout most of the paper (up to a few exceptions) we will omit to indicate it in the notation for the functions and sets related to (2.4). Thus, we will simply write
$\Phi_{\sigma}(u),\quad\phi(\sigma),\quad J_{\sigma},\quad\mathsf{d}^{\pm}_{\sigma}\,.$
@@
$\text{there exists a \emph{measurable} selection }(0,\sigma_{*})\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ (2.8)
To show this, let us consider the multivalued mapping
$\bm{J}:(0,\infty)\rightrightarrows X,\qquad\bm{J}(\sigma):=J_{\sigma}.$ (2.9)
It is easy to check that $\bm{J}$ is upper semicontinuous from $\mathbb{R}$} to $(M,{\mathcal{D}})$ in that it fulfills for every $(\sigma_{n})_{n},\,\sigma\in(0,\infty)$
$\sigma_{n}\to\sigma\ \Longrightarrow\ \mathop{\mathrm{Ls}}_{n\to\infty}\bm{J}(\sigma_{n})\subset\bm{J}(\sigma)\,$, (2.10a)
where the Kuratowski upper limit (cf., e.g., [AmT04, Def. 4.4.13]) of the sequence of closed sets $(\bm{J}(\sigma_{n}))_{n}$ is defined by
$u\in\mathop{\mathrm{Ls}}_{n\to\infty}\bm{J}(\sigma_{n})\ \Longleftrightarrow\ \exists\,(\sigma_{n_{k}})_{k},\,(u_{k})_{k}\text{ with }\begin{cases}\displaystyle u_{k}\in\bm{J}(\sigma_{n_{k}})\text{ for all }k\in{\mathbb{N}}\,,\\ \lim_{k\to\infty}{\mathcal{D}}(u,u_{k})=0\,.\end{cases}$ (2.10b)
Since $\bm{J}$ is upper semicontinuous, its graph is a Borel subset of $(0,\infty)\times X$, and the von Neumann-Aumann selection theorem [CaV77, Thm. 3.22] applies, yielding (2.8).
@@
Definition 2.7 (De Giorgi variational interpolant)
@@
The cornerstone of the proof of the convergence, as $\tau\downarrow 0$, of (a subsequence of) the sequence $(\widetilde{u}_{\tau})_{\tau}$ to a curve of maximal slope is the discrete estimate obtained by applying the De Giorgi estimate (1.4a) to the interpolant $\widetilde{u}_{\tau}$. The next section revolves around the validity of (1.4a) as an equality.
2.1 Metric energy identity and the De Giorgi estimate: statements and examples